OpenStudy (anonymous):

6y - 15 - 8y = 24 - 5

OpenStudy (pfenn1):

\[6y - 15 - 8y = 24 - 5\]rearrange and solve for y\[-2y=34\]\[y=-17\]

OpenStudy (anonymous):

6y - 15 - 8y = 24 - 5 -2y = 15 + 19 -2y = 34 divide equation by -2. y = 17

OpenStudy (anonymous):

thank you sooooo much

OpenStudy (pfenn1):

you're welcome. @vgwapo23, didn't you drop the negative sign somewhere?

OpenStudy (anonymous):

wat do you mean?

OpenStudy (anonymous):

yeah cause you divide it by -2.